What is the product of 20%, 48, and 50?
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
What is the product of 20%, 48, and 50?

A SOLUTION:
Before you multiply a percent, convert it to the equivalent decimal.

20 % becomes 0.20

Now multiply in two steps

0.20 * 48 * 50 = _____

0.20 * 48 = 9.6

9.6 * 50 = 480

If you want to know more about multiply decimal numbers and whole numbers together,
here's a good link. There are some practice problems at the bottom of the page with 
explanations worked out.

http://www.mathsisfun.com/multiplying-decimals.html


Hope this helps! Feel free to email if you have any questions.

Mrs. Figgy
math.in.the.vortex@gmail.com